Sol Summary – November 17th

Summary Title: First harvest
Author’s name: Izabela Shopova
Mission Status: Nominal
Sol Activity Summary: Two EVAs – collecting samples from White Rock Canyon and on-foot orientation expedition near the Hub. In the morning we completed the second session of the First Aid training, the crew engineer repaired the kitchen sink drain, and in the afternoon the crew scientist planted mushrooms and harvested the first batch of microgreens for our dinner salad.
Look Ahead Plan: Tomorrow we have planned two EVAs back-to-back in the afternoon and will continue working on on-going projects.
Anomalies in work: None
Weather: Clear, sunny and warm for the season.
Crew Physical Status: Nominal.
EVA: Two successful EVAs: Morning EVA with Opportunity and Perseverance to White Rock Canyon; Afternoon EVA on-foot in the area of the Hub.
Reports to be filed:
Commander’s report
EVA report
EVA request
Operations report
Sol Summary report
Support Requested: None
Thank you, Mission support.

Operations Report – November 17th

SOL: 4
Name of person filing report: Jas Purewal
Non-nominal systems: Pipe under sink was fixed using plumbing adhesive. Propane tank was 35% and refilled. Now full.
Notes on non-nominal systems: NA
Spirit rover used: No
Hours: (before EVA) NA
Beginning charge: (Before EVA) NA
Ending charge: (On return from EVA, before recharging) NA
Currently charging: No
Opportunity rover used: Yes
Hours: 103.3
Beginning charge: 100%
Ending charge: 54%
Currently charging: Yes
Curiosity rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: No
Perseverance rover used: Yes
Hours: 248.9
Beginning charge: 100%
Ending charge: 41%
Currently charging: Yes
General notes on rovers: All working fine
Summary of Hab operations: put notes here
WATER USE: 10 gallons
Water (static tank): 400 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off) off
Toilet tank emptied: No
Summary of internet: Good
Summary of suits and radios: optimal
Summary of GreenHab operations: None
WATER USE: 5.7 gallons
Heater: On
Supplemental light: Off
Harvest: Radish – 70.87g; Assorted Kale – 48.19g
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, Off
Summary of RAM operations: (Tools used, work done) Plumbing adhesive and tape removed from RAM and currently in hab.
Summary of any observatory issues: NA
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support: Still waiting on more trash bags please for toilet bins and for general bins. Pillows please.

GreenHab Report – November 17th

Sol 4
GreenHab Officer: Caitlyn Hubric
Environmental control: (heater, fan, door open?) heater on, door closed, 20% humidity
Average temperatures: 43.3ºC/13.5ºC
Hours of supplemental light: N/A
Daily water usage for crops: 20 L ≈ 5.28 gal
Daily water usage for research and/or other purposes: 1.6 L ≈ 0.42 gal
Water in Blue Tank ~282.3 gallons
Time(s) of watering for crops: twice, morning and night, 9am-6pm
Changes to crops:
-Added wooden sticks with sugar snap peas for support earlier today. When I checked tonight, a few of the plants had already begun to wrap around them.
-I tasted and used a plant identification app and confirmed that the unlabeled bin of microgreens was assorted kale.
-I harvested only half of the radish microgreens so we can consume more in a few days. The assorted kale I harvested came from the gray unlabeled bin located to the right of the metal shelving containing 3 rows of pots.
-Please see photos below of the harvest.
Harvest: (include which crop and mass in grams)
-Radish microgreens – 70.87 g
-Assorted kale – 48.19 g
Support/supplies needed: N/A

EVA Report – November 17th

EVA # 4
Author: Izabela Shopova
Purpose of EVA: Return to collect additional biological samples. Additional exploration of sites prospected during EVA #3 on Sol #3.
Start time: 10 AM
End time: 12 PM
Narrative: The crew went to the White Rock Canyon to collect more samples. Comms with CapCom were lost on the way and crew changed to channel 2 for communication between themselves. We set off on foot exploring and noted petrified wood and many types of rock in the area. Crew collected multiple lichen and regolit samples, checked two rock overhangs for life and took photos of the area. Found the expedition route ended back at the vehicles, where the two drivers made a tri-point u-turn and departed on time. Midway we reestablished radio contact with CapCom and returned to the base right on time. Commander Jenni checked the propane level before joining the rest of the crew in the airlock.
Destination: White Rock Canyon
Coordinates (use UTM WGS 84): 520000, 4248000
Participants: CDR, COMM, BIO
Road(s) and routes per MDRS Map: Cow Dung Rd (0110) to White Rock Canyon on Rover
Mode of travel: Rover, Walking (Oppy & Percy)

EVA # 5
Author: Judi Marcos
Purpose of EVA: Collect rocks for use with mechanical anchor prototype. (Note: Later engineering EVAs (Week 2) will use prototype in the field.)
Start time: 2:30 PM
End time: 4:00 PM
Narrative: Second crew decided to do a walking EVA on road 1103, wrapping around on to Hab Ridge. At the beginning of the EVA, the crew did a walk around of the hab for inspection. The hab continues to be in good condition. While on the EVA, we were able to see an amazing view of the land around after hiking to higher ground. We also retrieved a variety of rocks regarding shape and size. These sample rocks will be weighted to test carry capacity.
Destination: Hab Ridge Road North
Coordinates (use UTM WGS 84): 518000, 4251000 to 518000, 4250000
Participants: XO, ENGR, HSO
Road(s) and routes per MDRS Map: Walking north from Hab via Sagan Street to Hab Ridge Road North (518000, 4251000), then south along Hab Ridge Road (518000, 4250000). Returning via the same path.
Mode of travel: Walking

Commander Report – November 17th

Sol: 4
Summary Title: First Harvest
Author’s name: Dr. Jenni Hesterman, Commander
Mission Status: Nominal
Commander Report: The crew was up early and enjoyed breakfast burritos and fresh yogurt for breakfast. We then discussed our goals for the day and updated the EVA checklist with lessons learned.

After additional first aid training on the Oculus VR headset, 3 crew members took rovers out for a two-hour EVA to collect lichen, soil and rock samples from White Rock Canyon. They explored the dry riverbed, searching for fossils and plant life.

After a lunch of miso soup and dried seaweed, the other half of the crew set out on foot to Hab Ridge Road for an orientation expedition near the station. They also completed an inspection of the exterior of the habitat.

Additional activities included the crew engineer repairing the kitchen sink drain. Also, the crew biologist inoculated mushroom spores into substrate, which included coffee grounds and other food waste generated by the crew. She also harvested the first batch of microgreens for a dinner salad, which was served with chili and fresh bread.

Although we all just met for the first time a few days ago, the crew is already a high performing team. We attribute this to the months spent mission planning and getting to know each other on weekly video calls.

We wake up each day thankful for this incredible experience, endeavoring to maximize our time here on Mars.

Ad astra!

Copyright © The Mars Society. All rights reserved. | Main Site